# Why do Machine Learning Notebooks Crash?

### v1

@@ 3.3 Sampling
Our study requires a manual labeling and review process to analyze ML notebook crashes. Since investigating the entire population is infeasible, we select a meaningful sample of crashes for manual analysis similarly to De Santana et al. [8].
@@
We then cluster the preprocessed error values using Jaccard similarity [25] which measures overlap between sets. This method is ideal for grouping error values with similar patterns, as shown in previous studies [26, 27]. An example of clustering results using Jaccard similarity is presented in Table III. We chose a threshold of 0.7 to perform Jaccard similarity based on preliminary experiments. We first cluster all crashes (both GitHub and Kaggle) to get a comprehensive view of common patterns, then calculate clusters for GitHub and Kaggle datasets separately to identify platform-specific differences in crash behaviors. Each pair of error values is compared, and if their similarity exceeds the threshold, they are connected. Crashes are grouped into the same cluster if they can be reached through any connected crashes, ensuring reachability within the cluster. We obtain a total of 7,433 clusters. Separate statistics for each platform reveal 7,222 clusters for GitHub and 640 for Kaggle.
TABLE III: Jaccard similarity result of two error values.
Error value (original) Error value (processed) Jaccard similarity
name ’x’ is not defined name is not defined 1.0 (the same)
name ’head’ is not defined name is not defined
We further visualize the clusters in relation to exception types. We observe that multiple exception types have similar error values as Table IV shows. This is likely because Python exceptions can be customized and raised by users or library developers under various names or reasons. As a result, exceptions with different names may represent similar issues. This finding motivates our approach of clustering based on error values, rather than using exception types directly.
TABLE IV: Top 5 clusters across GitHub and Kaggle datasets. The first column shows cluster sizes (CS) with the number of exception types (#ET) per cluster. The second column lists the top 3 exception types and an example error value. Python built-in exception types are explained, while customized exceptions are marked as “self-defined.”
CS (#ET) Top 3 ET (exception types) Example error value
25070 (3) NameError(25050): local/global name not found UndefinedVariableError(19): self-defined RemoteError(1): self-defined
name ’sc’ is not defined
6107 (6) AttributeError(6072): attribute reference/assignment fails ModuleAttributeError(18): self-defined TypeError(13): operation/function apply to wrong type object
’numpy.ndarray’ object has no attribute ’sample’
5136 (3) ModuleNotFoundError(4568): module could not be located ImportError(567): cannot load a module DependencyNotinstalled(1): self-defined
No module named ’src.utils.compute_overlap’
4620 (83) KeyError(3911): a mapping key is not found SystemExit(349): raised by the sys.exit() ValueError(48): argument with right type but wrong value
’Hero Over’
3410 (16) FileNotFoundError(3230): file/directory not exist IOError(79): aliases of OSError OSError(28): system-related error,  e.g., ’disk full’
No such file or directory: ’GloVe/glove.840B.300d.txt’
@@
RQ1 exception type Python built-in exception types (e.g., NameError, ValueError, TypeError, KeyError) These are exceptions that occur during the execution of Python code, typically caused by issues like undefined variables, incorrect data types, or illegal key errors. Some examples can be seen in Table IV, more details can be found in the Python documentation.
others (e.g., APIArgumentError, EnvironmentError, ResourceError) These consist of library customized exceptions. For example, APIArgumentError arise when argument values fail to meet expected constraints of an API, EnvironmentError include import exceptions and system configuration issues, ResourceError contain exceptions such as out of memory or out of disk space.
root cause data confusion (DATA) Errors occur when the expected data structure (e.g., shapes, value types, feature names) does not match the actual input, or previously saved ML models (considered as external data) fails to align with the model defined in the code (e.g., model structure).
API misuse (API) Using an API incorrectly, e.g., providing wrong arguments or applying an API improperly.
notebook-specific issues (NB) Errors that stem from the unique execution behavior of notebooks, e.g., out-of-order execution, and unresolved errors in prior cells (previous cell error).
implementation error (IMPL) Faults in the coding logic, e.g., incorrect implementation or missing appropriate conditions.
library change (LIB) Errors caused by updates to an API or libraries, making current code incompatible.
environment issues (ENV) Issues due to incorrect environment setting, e.g., missing/incompatible libraries, file path errors.
intentional (INTENT) Errors deliberately introduced by users, often for testing or education purposes.
insufficient resources (RSC) Crashes due to resource constraints, like running out of memory or disk space.
@@
RQ3 ML pipeline environment setup (ENVS) Errors occurring when setting up the environment, e.g., missing modules or configuration issues.
@@
model training (MTRAIN) Errors that arise during model training, including hyperparameter tuning and validation.
@@
no ML pipeline (NONE) Errors unrelated to ML pipeline (e.g., notebooks for tutorials, education, or simulations).
@@
We estimate sample sizes using random sampling without replacement [28, 29]. We calculate sample sizes and analyze sampled crashes separately for GitHub and Kaggle because they may exhibit different behaviors. The desired sample sizes are determined with a 95% confidence interval (CI) and a 5% margin of error (MoE), aiming for results that are 95% accurate within a 5% margin of error from the entire population. This results in sample sizes of 360 for GitHub and 332 for Kaggle crashes.
@@ 3.4 Crash analysis
We manually refine the raw exception types for two reasons. First, we distinguish between Python’s built-in exceptions and those defined by third-party libraries. Our analysis reveals that 7.21% of crashes on GitHub and 7.31% on Kaggle have names differing from Python’s built-in types, indicating these exceptions are likely created by third-party libraries or users. By refining exception types, we can better identify library-specific exceptions. Moreover, as shown in Table IV, different exception types may exhibit similar patterns, while the same exception type may represent different triggers. Users can define and raise exception types freely, which may not always accurately represent the true trigger of a crash. Therefore, manually refining these types allows us to enhance our analysis of the crashes.
@@
We base our initial root cause categories on prior studies [7, 11, 8]. We also explore how notebook semantics, such as out-of-order execution and unresolved errors in previous cells, contribute to crashes. By categorizing crashes by root causes, we capture a broad range of crash sources, from misunderstanding of input data or APIs to environment-specific and notebook-specific issues, obtaining a comprehensive understanding of why the crashes happen.
@@
This dimension classifies crashes based on whether they occur in code related to ML libraries. Morovati et al. [12] compare ML bugs and non-ML bugs by identifying ML bugs as all bugs within ML components of DL systems. In contrast, our study focuses on the ML components themselves, aiming to evaluate how often crashes stem from direct interactions with ML libraries. We define ML bugs as crashes that occur in code using ML libraries or objects derived from them. By distinguishing ML bugs and general Python bugs, we can better understand the challenges unique to ML development, including the use of ML APIs, model handling, and data processing tasks.
@@
In this dimension, we categorize crashes based on where they occur during the ML pipeline to identify critical stages. Building on the framework from Islam et al. [7] and a pilot study of our dataset, we define key stages including environment setup, data preparation, data visualization, model construction, model training, and evaluation/prediction. This dimension is crucial for analyzing where in the ML pipeline crashes are more frequent and potentially harder to resolve.
@@ 3.5 Manual labeling process
In the labeling stage, we categorize crashes across the predefined dimensions introduced in Section 3.4. We adopt grounded theory methods [30], incorporating first-cycle and second-cycle coding phases [31]. We begin with an initial set of categories for each dimension, derived from prior studies [11, 7, 12] and our pilot study, and iteratively refines them as needed to maintain classification accuracy.
During the first-cycle coding, three evaluators independently label an initial set of 50 crashes from the sampled dataset. Following this, the evaluators meet to synchronize their findings, discuss the results, and establish uniform labeling principles. This collaborative discussion promotes consistency and mitigates potential bias early in the process.
@@
To further improve consistency, we organize crashes from both GitHub and Kaggle datasets by clusters, which groups similar crashes together for labeling and review. Each crash is categorized based on the predefined dimensions, with new categories introduced only when necessary. Additionally, evaluators reference all notebooks containing the sampled crashes to access relevant crash contexts. This structured approach, grounded in an open coding scheme [30], allows category expansion while minimizing bias and enhancing consistency in our classifications.
During the second-cycle coding, the remaining crashes are labeled independently. Regular meetings are held to discuss and reach consensus on any disagreements, promoting consistent standards across the entire dataset.
In the reviewing stage, a separate evaluator conducts reviewing to validate the labeling results. In cases of disagreement, discussions take place to resolve conflicts, and if needed, all three evaluators come together to reach a final decision. The change rates before and after reviewing are shown in Table VI. The process sometimes results in updates to earlier categories, particularly as the classification standards are refined during review. For example, many changes are made to categorize crashes in ML pipeline based on the stage where the code is written, rather than by its functionality, which leads to a high rate of reclassification. On the other hand, the high change rate of root cause reflects the challenges of accurately identifying the root causes of the crashes.
@@
TABLE VI: Changes of labeled crashes after reviewing.
RQ Dimension Changed total Percentage
RQ1 exception type 37 14.51%
root cause 81 31.76%
RQ2 ML/Python bug 25 9.80%
RQ3 ML pipeline 82 32.16%
After manually labeling and reviewing, we find four irrelevant crashes, such as cells containing random text like “end,” or exceptions raised to print comments like “Done!” These do not represent actual crashes and are excluded from our analysis, resulting in 389 labeled crashes from GitHub and 353 from Kaggle. Despite this reduction, the sample sizes remain statistically significant with 95% confident level.
@@ 4 Results
Next, we present the results to answer each research question. To investigate correlations between two distributions, we use the $\chi^{2}$ test to assess whether there is a significant difference, with a null hypothesis stating no difference. We adopt a 95% confidence level, meaning that a $p$-value less than 0.05 indicates a significant difference. For distributions with small sample sizes ($<5$), we apply Barnard’s exact test [32], using Bonferroni correction [33] when the categories are non-binary.
4.1 RQ1: What are the predominant exception types and root causes of crashes in ML notebooks?
(a) Exception types
@@ 4.1 RQ1: What are the predominant exception types and root causes of crashes in ML notebooks?
Crashes in ML notebooks stem from various types of exceptions. We refine and enhance the raw exception types typically shown in the crash information. Additionally, we complement these classes by identifying their underlying root causes and discussing how they relate to exception types. The results are visualized in Figure 3 and Figure 4.
Exception types
3(a) highlights the top 10 most frequent exception types. The most common exception type is NameError, accounting for 30% of cases (GitHub: 31%, Kaggle: 28%), followed by ValueError at 12% (GitHub: 11%, Kaggle: 12%), TypeError at 11% (GitHub: 10%, Kaggle: 12%), IOError at 9% (GitHub: 8%, Kaggle: 11%), and APIArgumentError also at 9% (GitHub: 10%, Kaggle: 8%).
The $\chi^{2}$ test shows no significant difference in exception type distributions between GitHub and Kaggle ($p>0.05$).
Finding 1.1: The most common exception types of ML notebook crashes are NameError (30%), ValueError (12%), and TypeError (11%).
@@
3(b) visualizes the distribution of root causes. The most common ones are API misuse (19% overall; GitHub: 18%, Kaggle: 21%), notebook-specific issues (19% overall; GitHub: 20%, Kaggle: 18%), data confusion (18% overall; GitHub: 17%, Kaggle: 19%), environment issues (16% overall; GitHub: 17%, Kaggle: 16%), and implementation errors (15% overall; GitHub: 15%, Kaggle: 15%).
Finding 1.2: The main root causes of ML notebook crashes include API misuse (19%), notebook-specific issues (19%), data confusion (18%), environment problems (16%), and implementation errors (15%).
Notebook-specific issues are a major cause of crashes, ranking first on GitHub and third on Kaggle. The majority of these crashes result from out-of-order execution (GitHub: 64%, Kaggle: 75%), where cells are executed in an incorrect sequence. This occurs when users run notebook cells without considering dependencies. Another common cause is previous cell error (GitHub: 36%, Kaggle: 25%), where users encounter crashes in earlier cells but continue executing dependent cells without resolving the issues, resulting in additional crashes.
@@
Finding 1.3: Notebook-specific issues are the leading cause of crashes on GitHub ML notebooks and the third most common on Kaggle, with out-of-order execution being the most frequent (GitHub: 64%, Kaggle: 75%) problem.
Root causes of exception types
We merge both sample sets (GitHub and Kaggle) and plot the co-occurrence of exception types and their root causes in 4(a).
@@
Interestingly, we observe that root causes like API misuse, data confusion, and implementation errors contribute to a wide range of exceptions in ML notebooks. Notably, API misuse alone is responsible for at least 20% of crashes across multiple exception types, including ValueError (28%), TypeError (32%), APIArgumentError (63%), AttributeError (41%), KeyError (20%), and IndexError (27%). Meanwhile, data confusion contributes to over 30% of ValueError (43%), TypeError (35%), KeyError (55%), and IndexError (54%). These findings indicate that developers need to be attentive to the proper use of APIs and ensure that their data inputs are correctly formatted.
@@
Finding 1.4: API misuse and data confusion are the primary causes of a broad range of exception types.
(a) Co-occurrence of exception types and root causes.
(b) The percentage point (pp) difference, such that GitHub has $x$ pp more/fewer instances than Kaggle.
@@
Figure 4: Overview of the co-occurrence of crash exception types and root causes (see Table V for full names), and the difference in co-occurrence between Kaggle and GitHub.
Certain exception types are dominantly linked to a specific root cause. For instance, IOError is connected to environment issues, APIArgumentError to API misuse, and KeyError and IndexError to data confusion. However, exceptions like ValueError, TypeError, and AttributeError have a wider range of root causes, indicating that these crashes may arise from more complex interactions in the code. Specifically, ValueError and TypeError can arise from data confusion (43%, 35%), API misuse (28%, 32%), and implementation errors (12%, 16%). AttributeError is primarily caused by API misuse (41%), implementation errors (20%), and data confusion (16%).
Moreover, we find that 56% of NameError exceptions stem from notebook-specific issues which also account for 20% of KeyError occurrences. This highlights how the unique notebook semantics can introduce distinct challenges that manifest in particular exception types.
Finding 1.5: ValueError, TypeError, and AttributeError arise from diverse root causes. NameError is mainly due to notebook-specific issues.
4(b) highlights the absolute percentage differences in co-occurrence between GitHub and Kaggle notebook crashes. Barnard’s exact tests with Bonferroni correction show no significant difference in root cause distributions between GitHub and Kaggle for all exception types.
However, some differences appear in the sample. Data confusion and API misuse both cause TypeError 1.6 times more commonly on Kaggle than on GitHub (a percentage points (pp) difference of 16). On GitHub, these are 4.5 and 1.8 times more commonly caused by intentional bugs and implementation errors (16 and 9 pp difference, respectively). Moreover, KeyError crashes are 1.5 times more commonly caused by data confusion on GitHub (23 pp difference), whereas notebook-specific issues are 3 times more frequent on Kaggle (18 pp difference).
@@
4.2 RQ2: To what extent are crashes in ML notebooks caused by ML bugs compared with general Python bugs?

### v2

@@ 3.3 Sampling
Our study requires a manual labeling and review process to analyze ML notebook crashes. Since investigating the entire population is infeasible, we select a meaningful sample of crashes for manual analysis.
@@
We then cluster the preprocessed error values using Jaccard similarity [30] which measures overlap between sets. This method is ideal for grouping error values with similar patterns, as shown in previous studies [31, 32]. An example of clustering results using Jaccard similarity is presented in Table III. We choose a 0.7 threshold for Jaccard similarity, based on its use in a previous study [32] in the context of code clone detection. We also test smaller and larger thresholds (ranging from 0.5 to 0.9 in 0.1 intervals). After evaluating 50 clusters formed at each threshold, we find that the smaller thresholds lead to clusters with mixed error messages, while the larger ones create too many similar clusters. The 0.7 threshold provides a coherent and meaningful grouping, making it the optimal choice for our analysis. We first cluster all crashes (both GitHub and Kaggle) to get a comprehensive view of common patterns, then calculate clusters for GitHub and Kaggle datasets separately to identify platform-specific differences in crash behaviors. Each pair of error values is compared, and if their similarity exceeds the threshold, they are connected. Crashes are grouped into the same cluster if they can be reached through any connected crashes, ensuring reachability within the cluster. We obtain a total of 7,433 clusters. Separate statistics for each platform reveal 7,222 clusters for GitHub and 640 for Kaggle.
TABLE III: Jaccard similarity result (JS) of two error values.
Error value (original) Error value to compare JS
Input 0 of layer ”sequential_1” is incompatible with the layer: expected shape=(None, 32), found shape=(None, 38) input of layer is incompatible with the expected shape none found 0.79
Input 0 of layer sequential_1 is incompatible with the layer: expected ndim=3, found ndim=2. Full shape received: (None, None) input of layer is incompatible with the expected ndim found full shape received none
TABLE IV: Different clusters within Value Error exception.
Cluster Example error value Crash type
2255 operands could not be broadcast together with shapes (25,200) (25,50) Unsupported broadcast
559 Classification metrics can’t handle a mix of binary and continuous targets Data value violation
720 Input 0 of layer ”sequential” is incompatible with the layer: expected shape=(None, 4), found shape=(None, 128) Tensor shape mismatch
6862 cannot reshape array of size 0 into shape (2,2)
6473 shapes (25,) and (26,1) not aligned: 25 (dim 0) != 26 (dim 0)
@@
RQ1 Crash type Variable Not Found (VNF) A referenced variable is not defined or accessible. A refined type of Name Error.
Invalid Argument (IARG) A function receives an argument with an incorrect type, format, or value, or has missing or misordered arguments, violating API constraints.
IO Error (IO) A crash occurs when reading from or writing to a file, e.g., file/path not found, no permission to the file, invalid file type, or unable to parse or decode the file.
Module Not Found (MODULE) A required library or module is missing or not installed. A refined type of Name Error.
Tensor Shape Mismatch (TSHAPE) Tensor shapes are incompatible for the intended operation. A refined type of Value Error.
Data Value Violation (DVIOL) Dataset contains unexpected or invalid values that violate API constraints or user assumptions. A refined type of Value Error.
Out of Memory (OOM) The available memory of GPU/CPU is insufficient to execute the operation.
Environment Error (ENV) Issues related to system setup or import failures.
Request Error (RERR) A failure occurs when making an external API or network request.
Unsupported Broadcast (BCAST) Broadcasting fails due to incompatible array or tensor shapes. A refined type of Value Error.
Feature Name Mismatch (FNAME) Feature names do not match expected names or formats in ML model structure or dataset. A refined type of Value Error.
Model Initialization Error (INIT) Crash during model creation due to incorrect parameters or architecture.
Python built-in exception types (e.g., Attribute Error(ATTR), Key Error(KEY), Value Error(VALUE), Type Error(TYPE), Name Error(NAME), Index Error(INDEX), Runtime Error(RUNTIME), System Error, Zero Division Error) These are exceptions that occur during the execution of Python code, typically caused by issues like attribute access errors, illegal dictionary keys, invalid values, type mismatches, undefined function/class names, out-of-bounds indexing, execution failures, system-related issues, division by zero, and so on. More details can be found in the Python documentation.
Root cause API misuse (API) Incorrect usage of an API, e.g., passing invalid arguments or calling functions improperly.
Notebook-specific issue (NB) Errors that stem from the unique execution behavior of notebooks, e.g., out-of-order execution, and unresolved errors in prior cells (previous cell error).
Data confusion (DATA) Misinterpretation of data shapes, structure, or values, leading to unexpected errors.
Implementation error (IMPL) Mistakes in code logic, e.g., incorrect variable usage, function calls, or algorithm design.
Library issue (LIB) Errors originating from external libraries, including bugs or deprecated APIs.
Environment issue (ENV) Issues due to incorrect environment setting, e.g., missing/incompatible dependencies, file path errors, or external control.
ML model Mismatch (MODEL) A mismatch between a saved ML model and the expected model definition in the code.
Insufficient resource (RSC) Crashes due to resource constraints, e.g., running out of memory or disk space.
@@
Library cause Specific ML libraries The crash is caused by using one/more ML libraries (see Table II).
None The crash is neither ML-related bug nor caused by directly using any ML libraries.
RQ3 ML pipeline Environment setup (ENVS) Errors related to environment setup, e.g., missing modules or configuration issues.
@@
Model training (TRAIN) Errors that arise during model training, including hyperparameter tuning and validation.
@@
We further visualize the clusters in relation to exception types. We observe that one exception type may have different patterns that lead to different crash types as Table IV shows. This is likely because Python exceptions can be customized and raised by users or library developers under various names or reasons. This finding motivates our approach of clustering based on error values, rather than using exception types directly.
@@
We estimate sample sizes using random sampling without replacement [33, 34]. We calculate sample sizes and analyze sampled crashes separately for GitHub and Kaggle because they may exhibit different behaviors. The desired sample sizes are determined with a 95% confidence interval (CI) and a 5% margin of error (MoE). This results in sample sizes of 360 for GitHub and 332 for Kaggle crashes. The details of the sample size calculation process can be found in our GitHub repository [15].
@@ 3.4 Crash analysis
We manually refine the raw exception types for two reasons. First, we distinguish between Python’s built-in exceptions and those defined by library developers or users. Our analysis reveals that 7.21% of crashes on GitHub and 7.31% on Kaggle have names differing from Python’s built-in types. By refining exception types, we can better identify customized exceptions. Moreover, as shown in Table IV, one exception type may exhibit different patterns. Users can define and raise exception types freely, which may not always accurately represent the true trigger of a crash. Therefore, manually refining these types allows us to enhance our analysis of the crashes.
@@
We base our initial root cause categories on prior studies (API misuse [9, 13], data confusion [13], implementation error [13, 10], library issues [9], environment issue [10], insufficient resource [10]). We extend this initial set of categories based on further insights from our pre-study. We discover how notebook semantics, such as out-of-order execution and previous unresolved errors, contribute to crashes. Additionally, we identify ML model mismatches as a root cause, occurring when a previously saved model is incompatible with the ML model defined in the current code. By categorizing crashes by root causes, we capture a broad range of crash sources, from misunderstanding of data or APIs to environment and notebook-specific issues, obtaining a comprehensive understanding of why the crashes happen.
@@
This dimension classifies crashes based on whether they occur in code related to ML libraries. Morovati et al. [14] compare ML bugs and non-ML bugs by identifying ML bugs as all bugs within ML components of DL systems. In contrast, our study focuses on the ML components themselves, aiming to evaluate how often crashes stem from interactions with ML libraries. We define ML bugs as crashes that occur in code using ML libraries or objects derived from them. By distinguishing ML bugs and general Python bugs, we can better understand the challenges unique to ML development, including the use of ML APIs, model handling, and data processing tasks.
LIBRARY CAUSE
This dimension labels which ML library is directly responsible for the crash. In our ML/Python bug classification, a crash is considered an ML bug if the code involves any ML APIs or components, regardless of whether the issue directly stems from ML libraries. This dimension focuses specifically on crashes caused by ML libraries, meaning that the crash must occur as a direct use of the library. As such, this is a subset of ML bugs. This dimension can contribute to determine which ML libraries are more challenging to use.
@@
In this dimension, we categorize crashes based on where they occur during the ML pipeline to identify critical stages. We define key stages building on the framework from a prior study by Zhang et al. [13] (environment setup, data preparation, training, and evaluation/prediction stages) and a pre-study of our dataset (data visualization and model construction stages). This categorization does not imply that all notebooks adhere to a strict pipeline. For example, some only implement a subset of stages, while others contain the same stage multiple times (e.g., when they try and compare different preprocessing methods). Rather, it provides a structured way to analyze where crashes occur within a general ML workflow and where crashes in the ML pipeline are more frequent and potentially harder to resolve.
@@ 3.5 Manual labeling process
In the labeling stage, we categorize crashes across the predefined dimensions introduced in Section 3.4. We adopt grounded theory methods [35], incorporating first-cycle and second-cycle coding phases [36].
We first label crashes based on predefined initial categories for each dimension (see Section 3.4). We then iteratively refine these categories whenever new crash patterns emerge that do not fit within the existing ones, particularly for Crash type and Root cause. We do this by adding new categories and updating existing ones, which is common in grounded theory [35]. For example, the crash type Tensor Shape Mismatch was originally included in Value Error and later split into a separate category. This creates more consistent and well-separated categories.
@@
During the first-cycle coding, three evaluators independently label an initial set of 50 crashes from the sampled dataset. Following this, the evaluators meet to synchronize their findings, discuss the results, and establish uniform labeling principles. This collaborative discussion promotes consistency and mitigates potential bias early in the process. To further improve consistency, we organize crashes from both GitHub and Kaggle datasets by clusters, which groups similar crashes together for labeling and review. Additionally, evaluators reference all notebooks containing the sampled crashes to access relevant crash contexts. This structured approach, grounded in an open coding scheme [35], allows category expansion while minimizing bias and enhancing consistency in our classifications. During the second-cycle coding, the remaining crashes are labeled independently. Regular meetings are held to discuss and reach a consensus on any disagreements, promoting consistent standards across the entire dataset.
In the reviewing stage, a separate evaluator validates the labeling results. In cases of disagreement, discussions take place to resolve conflicts, and if needed, all three evaluators come together to reach a final decision.
@@
TABLE VI: Changes in the number and percentage of labeled crashes after reviewing the final of sampled crashes (746 crashes after resampling).
RQ Dimension Number Percentage
RQ1 Crash type 31 4.16%
Root cause 72 9.65%
RQ2 ML/Python bug 24 3.22%
Library cause 32 4.29%
RQ3 ML pipeline 64 8.58%
3.6 Resampling
After manually labeling and reviewing the data, we identify 123 cases in which the crashes do not fit into any stage of the ML pipeline. The majority of these cases originate from tutorial notebooks, educational materials, or physics simulations. Additionally, we find four irrelevant crashes (e.g., cells containing random text, exceptions raised to print comments). As these instances do not represent actual crashes in ML development, we exclude them from our analysis.
To maintain statistically significant sample sizes, we use a resampling approach similar to proportional-to-cluster-size sampling (Section 3.3), where replacements are drawn from the same clusters where exclusions occur. We then verify whether the replacements should be excluded as well. If such cases exist, we iteratively perform resampling until we reach the target sample sizes (GitHub: 390, Kaggle: 356). Afterwards, we manually label the 127 resampled crashes following the same process (see Section 3.5).
The change rates before and after reviewing are shown in Table VI. The reviewing process sometimes results in updates to earlier categories. For example, many changes are made to categorize crashes in ML pipeline based on the stage where the code is written, rather than by its functionality, which leads to a high rate of reclassification. On the other hand, the high change rate of Root cause reflects the challenges of accurately identifying the root causes of the crashes.
@@ 4 Results
Next, we present the results to answer each research question. To investigate correlations between two distributions, we use the $\chi^{2}$ test to assess whether there is a significant difference, with a null hypothesis stating no difference. We adopt a 95% confidence level, meaning that a $p$-value less than 0.05 indicates a significant difference. For distributions with small sample sizes ($<5$), we apply Barnard’s exact test [37], using Bonferroni correction [38] when the categories are non-binary. When specifying our results, we refer to GitHub as GH and Kaggle as KG.
4.1 RQ1: What are the predominant crash types and root causes of crashes in ML notebooks?
(a) Crash types (see Table V for the full names)
@@ 4.1 RQ1: What are the predominant crash types and root causes of crashes in ML notebooks?
Crashes in ML notebooks stem from various types of exceptions. We refine and enhance the raw exception types typically shown in the crash information and form crash types. We then complement these classes by identifying their underlying root causes and discussing their relations. The results are visualized in Figure 3 and Figure 4.
Crash types
3(a) highlights the most frequent crash types in both GitHub and Kaggle and crash types that are unique to ML. The most common crash type is Variable Not Found, accounting for 17.8% of cases (GH: 18.5%, KG: 17.1%), followed by Invalid argument at 9.8% (GH: 10.3%, KG: 9.3%), IO Error at 9.7% (GH: 9.0%, KG: 10.4%), Module Not Found at 8.0% (GH: 9.7%, KG: 6.2%), and Attribute Error at 7.0% (GH: 7.4%, KG: 6.5%).
Finding 1.1: The most common crash types of ML notebook crashes are Variable Not Found (17.8%), Invalid argument (9.8%), and IO Error (9.7%).
Moreover, crash types Tensor Shape Mismatch (5.5%; GH: 4.6%, KG: 6.5%), Data Value Violation (3.5%; GH: 3.9%, KG: 3.1%), Out of Memory (2.1%; GH: 1.3%, KG: 3.1%), Unsupported Broadcast (1.6%; GH: 1.3%, KG: 2.0%), Feature Name Mismatch (1.2%; GH: 1.5%, KG: 0.8%), and Model Initialization Error (0.7%; GH: 0.5%, KG: 0.8%) are ML-specific crash types.
The $\chi^{2}$ test shows no significant difference in crash type distributions between GitHub and Kaggle ($p>0.05$).
Finding 1.2: ML-specific crashes types (14.6%) include Tensor Shape Mismatch, Data Value Violation, Out of Memory, Unsupported Broadcast, Feature Name Mismatch, and Model Initialization Error.
@@
3(b) visualizes the distribution of root causes for GitHub and Kaggle separately. The most common ones are API misuse (20.9% overall; GH: 21.0%, KG: 20.8%), notebook-specific issues (19.4%; GH: 19.2%, KG: 19.7%), implementation errors (16.9%; GH: 17.4%, KG: 16.3%), environment issues (16.8%; GH: 18.0%, KG: 15.5%), and data confusion (16.4%; GH: 16.9%, KG: 15.7%).
Finding 1.3: The main root causes of ML notebook crashes include API misuse (20.9%), notebook-specific issues (19.4%), implementation errors (16.9%), environment problems (16.8%), and data confusion (16.4%).
Notebook-specific issues are a major cause of crashes, ranking second on both GitHub and Kaggle. The majority of these crashes result from out-of-order execution (GH: 60.0%, KG: 74.3%), where cells are executed in an incorrect sequence. This occurs when users run notebook cells without considering dependencies. Another common cause is previous cell error (GH: 40.0%, KG: 25.7%), where users encounter crashes in earlier cells but continue executing dependent cells without resolving the issues, resulting in additional crashes.
@@
Finding 1.4: Notebook-specific issues are the second major cause of crashes in ML notebooks, with out-of-order execution being the most frequent (GH: 60.0%, KG: 74.3%).
Root causes of crash types
We find no statistical difference in root causes between GitHub and Kaggle, hence we merge them and plot crash type and root cause co-occurrences in Figure 4.
@@
We observe that API misuse, implementation errors, and data confusion are the root causes of a wide range of crashes in ML notebooks. Notably, API misuse alone is responsible for at least 25% of crashes observed along multiple crash types, including Invalid Argument (66%), Attribute Error (44%), Type Error (42%), Value Error (35%), Data Value Violation (31%), and Tensor Shape Mismatch (29%). Meanwhile, data confusion contributes to over 30% of several crash types, such as Key Error (50%), Data Value Violation (50%), Unsupported Broadcast (50%), Index Error (45%), Value Error (42%), Tensor Shape Mismatch (34%), and Feature Name Mismatch (33%). These findings indicate that debugging is likely to be difficult in such cases. Developers need to be attentive to the proper use of APIs and ensure that their data inputs are correctly formatted.
@@
Finding 1.5: API misuse, implementation errors, and data confusion are main causes of a wide range of crash types.
Certain crash types are dominantly linked to one specific root cause. For instance, Out of Memory is always raised due to insufficient resources. IO Error, Environment Error, and Request Error are related to environment problems. Invalid Argument is mainly due to API misuse and data confusion. Name-related errors such as Variable Not Found, Name Error, and Module Not Found are mostly connected to notebook-specific issues. This highlights how notebook semantics can introduce unique challenges that manifest in particular crash types.
Finding 1.6: Name-related errors, in particular Variable Not Found, are mainly due to notebook-specific issues.
To assess the accuracy of our root cause analysis, we try to reproduce 30 crashes in Kaggle and successfully reproduce 20. We then resolve the crashes, finding that 18 ($90\%$) are caused by the root causes identified in this study. This high accuracy suggests that our approach is reliable.
@@
Figure 4: Overview of the co-occurrence of crash types and root causes (see Table V for full names).
@@
4.2 RQ2: To what extent are crashes in ML notebooks caused by ML bugs and by which ML libraries?